OpenStudy (surryyy):

Simplify the expression 7 - 2(2 +5) - 8

OpenStudy (anonymous):

7 - 2(7) - 8 = 7-14-8 = -15

imqwerty (imqwerty):

7-2(2+5)-8 7-8-2(2+5) 7-8-2(7) 7-8-14 -1-14 -15

OpenStudy (surryyy):

tysm. makes more sense now

imqwerty (imqwerty):

welcome :)